### v1

# captionUnknown document class

Abstract
Randomized benchmarking (RB) is the most commonly employed protocol for the characterization of unitary operations in quantum circuits due to its reasonable experimental requirements and robustness against state preparation and measurement (SPAM) errors. So far, the protocol has been limited to discrete or fermionic systems, whereas extensions to bosonic systems have been unclear for a long time due to challenges arising from the underlying infinite-dimensional Hilbert spaces. In this work, we close this gap and develop an RB protocol for passive Gaussian transformations, which we call bosonic passive RB. The protocol is based on the recently developed filtered RB framework [helsenGeneralFrameworkRandomized2020, heinrichRandomizedBenchmarkingRandom2023] and is designed to isolate the multitude of exponential decays arising for bosonic systems. We give explicit formulas and a Julia implementation for the necessary post-processing of the experimental data. We also analyze the sampling complexity of bosonic passive RB by deriving analytical expressions for the variance. They show a mild scaling with the number of modes, suggesting that bosonic passive RB is experimentally feasible for a moderate number of modes. We focus on experimental settings involving Fock states and particle number resolving measurements, but also discuss Gaussian settings, deriving first results for heterodyne measurements.
I Introduction
The characterization of quantum devices, and, in particular, of the involved unitary operations, is a fundamental task in quantum information processing [Kliesch2020TheoryOfQuantum, Eisert2020QuantumCertificationAnd]. While being a well-studied field for discrete variable systems, a similar standing for CV (CV) systems has not yet been achieved. In fact, there are only very few works apart from CV tomography and the first rigorous guarantees for the latter were proven only recently [wuQuantumEnhancedLearningContinuousVariable2023, meleLearningQuantumStates2024]. The required number of measurements is experimentally extremely challenging [lvovskyContinuousvariableOpticalQuantum2009, lopezProgressScalableTomography2010, schmiegelowSelectiveEfficientQuantum2011, benderskySelectiveEfficientQuantum2013, namikiSchmidtnumberBenchmarksContinuousvariable2016, altepeterAncillaAssistedQuantumProcess2003, baiTestOneTest2018] and as a result, full quantum tomography is often not feasible for CV systems. Moreover, tomographic protocols typically suffer from SPAM (SPAM) errors, as in the discrete variable case.
Bosonic quantum systems play a major role in the design of quantum computing platforms. Most prominently, this includes photonic quantum computing as a popular proposal for scalable quantum computers [knillSchemeEfficientQuantum2001, koashiProbabilisticManipulationEntangled2001, kokLinearOpticalQuantum2007, rudolph2016i, tanResurgenceLinearOptics2018, alexanderManufacturablePlatformPhotonic2024]. The biggest advantages of this model rely on the implementation of particle sources, detectors, and linear optical circuits on the same integrated chips [politiSilicaonSiliconWaveguideQuantum2008, sprengersWaveguideSuperconductingSinglephoton2011, silverstoneOnchipQuantumInterference2014, meanyLaserWrittenCircuits2015], and access to mixed schemes for quantum error correction [bourassaBlueprintScalablePhotonic2021]. Bosonic systems also offer interesting non-universal models of computation to test quantum supremacy, such as boson sampling [aaronsonComputationalComplexityLinear2010] and Gaussian boson sampling [hamiltonGaussianBosonSampling2017, chakhmakhchyanBosonSamplingGaussian2017], which have been implemented successfully in real experiments recently [zhongQuantumComputationalAdvantage2020, zhongPhaseProgrammableGaussianBoson2021, madsenQuantumComputationalAdvantage2022].
Most challenges in the characterization of bosonic systems can be attributed to the particularities of the infinite-dimensional Hilbert space [lvovskyContinuousvariableOpticalQuantum2009, sharmaCharacterizingPerformanceContinuousvariable2020]. For instance, protocols that rely on scrambling techniques via unitary designs are challenging to adapt, since Gaussian unitaries only form a unitary $1$-design [blume-kohoutCuriousNonexistenceGaussian2011, zhuangScramblingComplexityPhase2019], and unitary $2$-designs cannot exist for CV systems, unless rigged Hilbert spaces are considered [iosueContinuousvariableQuantumState2022]. Such issues constrain characterization protocols to very specific settings, where the implemented unitary operators (which may include non-Gaussian single-mode unitaries) are only benchmarked w.r.t. a restricted set of input states [fariasCertificationContinuousvariableGates2021].
For discrete variable systems, RB (RB) [EmeAliZyc05, Levi07EfficientErrorCharacterization, DanCleEme09, EmeSilMou07, KniLeiRei08, MagGamEme11, Magesan2012, proctor2017WhatRandomizedBenchmarking, wallman2018randomized, merkelRandomizedBenchmarkingConvolution2019] is the most widespread family of protocols for the estimation of average gate fidelities. Its popularity is due to its robustness against SPAM errors and its rather low demands on the measurement effort [HarHinFer19]. The standard RB protocol is as follows: To a fixed initial state, apply a sequence of Haar-random (Clifford) unitaries, followed by a final inversion gate cancelling the action of the entire sequence. Then, the success probability of restoring the initial state decays exponentially with the length of the sequence, and the decay rate is a proxy for the average gate fidelity of the gate set.
Recent efforts led to general guarantees for RB protocols with general finite or compact groups [francaApproximateRandomizedBenchmarking2018, helsenNewClassEfficient2019, helsenGeneralFrameworkRandomized2020, helsenEstimatingGatesetProperties2021, kongFrameworkRandomizedBenchmarking2021, heinrichRandomizedBenchmarkingRandom2023] instead of only unitary $2$-designs. Due to the non-existence of the latter, this generality is clearly important when considering RB of CV systems. However, there is still a fundamental problem: The RB signal generally consists of a linear combination of exponential decays in correspondence with the relevant irrep of the used group [francaApproximateRandomizedBenchmarking2018, helsenNewClassEfficient2019, helsenGeneralFrameworkRandomized2020]. Estimating these decay rates is already difficult in practice if more than a few irreps are involved [helsenGeneralFrameworkRandomized2020, heinrichRandomizedBenchmarkingRandom2023] and may become impossible in the CV setting.
Very recently, a bosonic randomized benchmarking protocol using random displacements has been put forward [valahu_benchmarking_2024]. Due to the Abelian character of the (projective) Heisenberg-Weyl group and the considered noise models, the complicated behaviour involving many decays is avoided in this setting, allowing for a simple analysis.
However, for more complicated – in particular non-Abelian – groups, one cannot hope for such a simple analysis. To resolve the problems coming with many irreps, filtered RB has been recently proposed [helsenGeneralFrameworkRandomized2020, heinrichRandomizedBenchmarkingRandom2023], building on insights from character RB [helsenNewClassEfficient2019]. This protocol omits the inversion gate and instead performs a suitable post-processing of the data. During the post-processing, contributions associated to individual irreps can be isolated in a SPAM-robust way, allowing to handle compact groups with many irreps.
In this work, we introduce a RB protocol for bosonic systems, bosonic passive RB (or passive RB for simplicity), for the benchmarking of passive (Gaussian) transformations. To this end, we build on previous work on filtered RB [heinrichRandomizedBenchmarkingRandom2023], and make the common assumption of gate-dependent, stationary, and Markovian noise. To circumvent problems arising from infinite dimensions, we consider experiments where the input state is a number state and PNR (PNR) measurements are performed, resembling boson sampling experiments [aaronsonComputationalComplexityLinear2010]. For practical reasons, we propose to use collision-free states with at most one particle per mode. We then show that the passive RB succesfully benchmarks the average quality of passive transformations on a fixed number of particles, thereby separating the effect of particle number-preserving noise from particle loss using different post-processing of the same experimental data. In general, we prove that there are as many decay rates as we have particles in the experiments – however it is sufficient to only estimate a constant number of them in practice. By representation-theoretic means, we derive explicit and concise formulas for the functions used in the post-processing and provide a Julia implementation. We simulate the passive RB protocol for a small number of modes, demonstrating that it works as intended. Finally, we analyze the sampling complexity of our protocol by computing the relevant variances. We evaluate the obtained expressions numerically and find a very mild scaling with the number of modes $m$ which may be logarithmic – a definite conclusion is however prevented by the small number of data points.
The effectively finite-dimensional setting is introduced for a good reason: Although arguably a simpler experimental setting, Gaussian input states and measurements result in infinitely many relevant irreps and thus decay rates, rendering a meaningful analysis impossible. As a first step, we also consider scenarios in which the input state is still a number state, but each mode is measured with a (balanced) heterodyne detector. In this cases, passive RB is similar to Gaussian boson sampling experiments [hamiltonGaussianBosonSampling2017, chakhmakhchyanBosonSamplingGaussian2017].
The remainder of this work is structured as follows. In Section II we highlight our main results: LABEL:sec:protocol spells out our passive RB protocol in full detail. In LABEL:sec:guarantees we analyze our protocol, discuss the role of input states, and prove rigorous guarantees including bounds on its sampling complexity. LABEL:sec:discussion is devoted to a general discussion of the passive RB protocol, open problems, and possible extensions. The technical details and proofs are deferred to the technical part, LABEL:sec:passive_RB_Fock. Finally, LABEL:sec:passive_rb_Gaussian contains an analysis of passive RB with balanced heterodyne measurements.
II Main results
II.1 Notation
We consider a bosonic system of $m\in\mathbb{N}_{\geq 2}$ modes described by annihilation and creation operators $a_{k}$ and $a_{k}^{\dagger}$ ($k=1,\dots,m$), respectively, satisfying the CCR
$[a_{k},a_{l}^{\dagger}]=\delta_{kl}\,,\quad[a_{k},a_{l}]=[a_{k}^{\dagger},a_{l}^{\dagger}]=0\,,\quad k,l=1,\dots,m\,,$
where $\delta_{k,l}$ is the Kronecker delta. These operators act on the Fock-Hilbert space $\mathcal{F}_{m}\coloneqq\bigoplus_{n=0}^{\infty}\mathcal{H}_{n}^{m}$, where $\mathcal{H}_{n}^{m}$ is the subspace of $n$ bosons distributed over $m$ modes of dimension $\dim\mathcal{H}_{n}^{m}=\binom{n+m-1}{n}$. It is spanned by the Fock or number states
$|\boldsymbol{n}\rangle\equiv|n_{1},\dots,n_{m}\rangle\coloneqq\prod_{k=1}^{m}\frac{1}{\sqrt{n_{k}!}}a_{k}^{\dagger n_{k}}|\mathbf{0}\rangle\,,$ (1)
where $|\boldsymbol{n}|\coloneqq\sum_{i=1}^{m}n_{i}=n$ and $|\mathbf{0}\rangle\equiv|0,\dots,0\rangle$ denotes the vacuum state of $m$ decoupled one-dimensional harmonic oscillators [adessoContinuousVariableQuantum2014]. We shall also consider coherent states, defined as $|\alpha\rangle=e^{-\lvert}$